OpenStudy (anonymous):

If the first term in an arithmetic sequence is –3 and the tenth term is 15, what is the common difference? A. d = 2 B. d = 3 C. d = 6 D. d = 12

OpenStudy (anonymous):

helllllllp

OpenStudy (anonymous):

ok... Hi!

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

? @ayanna1100

OpenStudy (anonymous):

okay then. can someone else please help I have limited time

OpenStudy (welshfella):

first term = a1 = -3 tenth term = a1 + 9d = 15 so 15 = -3 + 9d solve for d

OpenStudy (welshfella):

because nth term of an AS = a1 + (n-1)d

OpenStudy (anonymous):

ok so a. thanks!

OpenStudy (welshfella):

right yw